OpenStudy (anonymous):

PLEEAASSEEE HEELLPPP!! NEED STEPS! Match each equation with the correct solution.

OpenStudy (anonymous):

\[4^{p-1}=3^p\]

OpenStudy (anonymous):

here are the answer choices 0.5537 4.8188 0.4907 2.0145 2.4450 4.8362

OpenStudy (anonymous):

@arch1 can you help me?

OpenStudy (anonymous):

Take log(10) both the sides. 10 is the base..

OpenStudy (anonymous):

\[\large \log_{10}(4)^{p-1} = \log_{10} (3)^p\] Use power rule here: \[\huge \log_n(x)^a = a \cdot \log_n(x)\]

OpenStudy (anonymous):

\[(p-1) \cdot \log_{10}(2)^2 = p \cdot \log_{10}(3)\] Again use power rule here: \[(2p-2) \cdot \log_{10}(2) = p \cdot \log_{10}(3)\]

OpenStudy (anonymous):

Use: \[\log_{10}(2) = 0.3010\] \[\log_{10}(3) = 0.4771\] And now find p..